OpenStudy (faman39):

Please help!!!!!!!!!!! Evan is painting a stool that has a cylindrical seat and four rectangular prism legs. The seat has a diameter of 14 inches and a height of 4 inches. Each leg is 4 inches by 4 inches by 18 inches. How much area will Evan have to paint? Use 3.14 for pi . Show your work.

OpenStudy (faman39):

@JakeV8 can you able to help me with this?

OpenStudy (anonymous):

you have to break it all up into parts to find the total area. SEAT: The cylindrical seat has surface area of pi*radius^2 on the top side. The side of the cylinder seat has surface area 2*pi*r*height I am gonna guess and say Evan should paint the whole bottom of the stool seat too, even though the legs have to connect there. Unfortunately, your question is very vague about this... but let's assume you paint the bottom... that's another pi*radius^2

OpenStudy (anonymous):

so, the seat radius is half the diameter, and the diameter is 14, so radius is 7. The top and bottom areas are pi * radius^2, or 3.14 * 7 * 7 = 153.86 Together, that totals to 307.72 for the top and bottom. The side area is 2 * pi * radius * height = 2 * 3.14 * 7 * 4 = 175.84. So the seat total is the top + bottom + side = 307.72 + 175.84 = 483.56

OpenStudy (faman39):

Thank you so much @JakeV8 :) and i m sorry for trouble, i m really a worst person

OpenStudy (anonymous):

You can probably do the legs... they are boxes with 4 x 4 square tops and bottoms, and all four sides are 4 x 18. And don't forget, there are four legs. Again, we're making the assumption that Evan has to paint the entire leg...top, bottom, and sides... if the legs were already attached to the stool this is a bad assumption, since he wouldn't be able to paint the top surfaces. Also, if this is true, you would have to subtract off 4 times the 4x4 top surface from the bottom of the cylinder, again, since that part of the bottom wouldn't be reachable since the legs would be attached. It's too bad the question is so vague about that...

OpenStudy (faman39):

you know what? i m not good at anything.

OpenStudy (anonymous):

you don't have to be good... the bottoms of the legs are squares 4 inches on a side... area = 4 x 4 = 16 the sides are rectangles 4 inches across and 18 inches tall... area = 4 x 18 = 72 and each leg has 4 sides... so 4 x 72 and if you add the top of each leg, that's another 4 x 4 square, area = 16. So each full leg has area 16 + 4(72) + 16

OpenStudy (faman39):

Ok, i will try understand cuz it is a long process

OpenStudy (faman39):

hard to understand

OpenStudy (anonymous):

yes, this one is a pain... lots of steps... easy steps, but there are a lot of them.

OpenStudy (faman39):

Ok thank u much for helping and tolerating me :) i will try to understand this which u show me. thanks alot for help

OpenStudy (anonymous):

you can do area... Just make a list of all the many parts of this stool that you need to find area on... that way you don't forget one part. There should be a top, bottom, and side of the cylinder. There should be 4 legs, each with a top, bottom, and 4 sides.